If a 61 g sample of metal at 100oC is placed into a calorimeter containing 60.0 g of water at 18.0 oC, the temperature of the water increases to 22.0 °C.

Calculate the amount of heat lost by the metal sample in calories.
Answer #1

According to the first law of thermodynamics heat release equal to heat accepted by other body. Hence amount of heat is released by the metal will be equal to the heat accepted by water of calorimeter. So amount of heat the raise the temparature is H = m * s * T

here m, mass of water.

s, specific heat of water.

T , tempatature change. (22-18) = 4

So H = 60 * 1 * 4 = 240 cal, = 1.008 jule